# Designing a skilled soccer team for RoboCup: exploring Skill-Set-Primitives through reinforcement learning

### v1

@@ 7. Conversion
Reset
One robot from our team approaches the ball, aligns itself with it, and begins the approach until close enough. If it falls or gets overtaken by a teammate, the active player is switched, and the procedure is restarted.
Termination
Robot has fallen ($z<0.40$), timeout (8 s), ball was lost, ball cannot be seen, teammate is closer to ball
Episode
The opposing team executes 1 time step: the closest opponent to the ball is physically simulated and uses the Omnidirectional Walk to push the ball to our goal. The other 10 opponents adhere to a simplified locomotion model, use FC Portugal’s formation, and are represented as 2D points with velocity and acceleration. Our team executes an analogous behavior, except for the current closest player, which is the only player controlled by the policy being learned.
Reward
$r=\min(\text{ball\_speed},0.5)\times\max(0,\cos(\text{direction\_error}))^{7}$, where direction_error is the deviation, in radians, between the ball velocity vector and the target position vector generated by Push HL. The ball speed is limited to 0.5 m/s to prioritize ball possession and progression over high speed. The reward is updated every vision cycle.
@@
radar
160
Radial spatial segmentation where each segment indicates the presence of opponents and/or teammates (see Section 6.2.1)
Sagittal reflectiona
Sagittal reflectiona
mate
1
Closest teammate Euclidean distance
0
1
opp
1
Closest opponent Euclidean distance
0
1
targdir
1
Last HL target direction, relative to user-defined long-term goal direction
0
1
a The symmetry operation associated with this sagittal reflection depends on the radar’s sensor order
@@
1. Action
Target direction (1 value)
2. Map
The direction is mapped to the range [-90,90]
3. Filter
A filtered direction is set to a maximum variation of 30 deg per iteration
@@ 3. Filter
Reset
Wait for Push LL to reset
Termination
Push LL resets and, simultaneously, ball exits learning area ($x\in[-12,12],y\in[-9,9]$)
Episode
Push LL is executed 16 times prior to re-observing the environment and running the policy, at a frequency of 3.125 Hz.
Reward
@@ A.5 List of videos
Skill Video IDa Description
Sprint-Kick
Yy1yCM5hwZI
Sprint-Kick training environment
3MND8RVUPBQ
Successful Sprint-Kick Kickoffs in RoboCup 2021
udN2F3oXAec
Sprint-Kick slow motion
Omnidirectional Walk
dXzIuZlOFZc
Omnidirectional Walk training environment and demonstration
Short Kick
trn1KzXFhwc
Short Kick training environment
Long Kick
pu4g5wryfEs
Best kick goals in RoboCup 2022
QQXGgpcnVYE
Long Kick test to assess the final ball position
Dribble
8UED_Zl-nbQ
Dribble evolution, training environment and demonstration
Push
rGWN83FBdJ4
Push training environment
mhfQ0NAmlMI
Push highlights from Brazil Open 2022
a Access through https://youtu.be/{Video ID}

### v2

@@ 6. IK Apply inverse kinematics to obtain the target joint positions
Reset One robot from our team approaches the ball, aligns itself with it, and begins the approach until close enough. If it falls or gets overtaken by a teammate, the active player is switched, and the procedure is restarted.
Termination Robot has fallen ($z<0.40$), timeout (8 s), ball was lost, ball cannot be seen, teammate is closer to ball
Episode The opposing team executes 1 time step: the closest opponent to the ball is physically simulated and uses the Omnidirectional Walk to push the ball to our goal. The other 10 opponents adhere to a simplified locomotion model, use FC Portugal’s formation, and are represented as 2D points with velocity and acceleration. Our team executes an analogous behavior, except for the current closest player, which is the only player controlled by the policy being learned.
Reward $r=\min(\text{ball\_speed},0.5)\times\max(0,\cos(\text{direction\_error}))^{7}$, where direction_error is the deviation, in radians, between the ball velocity vector and the target position vector generated by Push HL. The ball speed is limited to 0.5 m/s to prioritize ball possession and progression over high speed. The reward is updated every vision cycle.
@@
radar 160 Radial spatial segmentation where each segment indicates the presence of opponents and/or teammates (see Section 6.2.1) Sagittal reflection [^1] The symmetry operation associated with this sagittal reflection depends on the radar’s sensor order Sagittal reflection [^1] The symmetry operation associated with this sagittal reflection depends on the radar’s sensor order
mate 1 Closest teammate Euclidean distance 0 1
opp 1 Closest opponent Euclidean distance 0 1
targdir 1 Last HL target direction, relative to user-defined long-term goal direction 0 1
@@
1. Action Target direction (1 value)
2. Map The direction is mapped to the range [-90,90]
3. Filter A filtered direction is set to a maximum variation of 30 deg per iteration
@@
Reset Wait for Push LL to reset
Termination Push LL resets and, simultaneously, ball exits learning area ($x\in[-12,12],y\in[-9,9]$)
Episode Push LL is executed 16 times prior to re-observing the environment and running the policy, at a frequency of 3.125 Hz.
@@ A.5 List of videos
Skill Video ID [^1] Access through https://youtu.be/{Video ID} Description
Sprint-Kick Yy1yCM5hwZI Sprint-Kick training environment
3MND8RVUPBQ Successful Sprint-Kick Kickoffs in RoboCup 2021
udN2F3oXAec Sprint-Kick slow motion
Omnidirectional Walk dXzIuZlOFZc Omnidirectional Walk training env. and demonstration
Short Kick trn1KzXFhwc Short Kick training environment
Long Kick pu4g5wryfEs Best kick goals in RoboCup 2022
QQXGgpcnVYE Long Kick test to assess the final ball position
Dribble 8UED_Zl-nbQ Dribble evolution, training environment and demonstration
Push rGWN83FBdJ4 Push training environment
mhfQ0NAmlMI Push highlights from Brazil Open 2022